# references.bib

### v1

@@ Abstract.
In this work, we develop a new compatible finite element formulation of the thermal shallow water equations that conserves energy and mathematical entropies given by buoyancy–related quadratic tracer variances. Our approach relies on restating the governing equations to enable discontinuous approximations of thermodynamic variables and a variational continuous time integration. A key novelty is the inclusion of centred and upwinded fluxes. The proposed semi-discrete system conserves discrete entropy for centred fluxes, monotonically damps entropy for upwinded fluxes, and conserves energy. The fully discrete scheme reflects entropy conservation at the continuous level. The ability of a new linearised Jacobian, which accounts for both centred and upwinded fluxes, to capture large variations in buoyancy and simulate thermally unstable flows for long periods of time is demonstrated for two different transient case studies. The first involves a thermogeostrophic instability where including upwinded fluxes is shown to suppress spurious oscillations while successfully conserving energy and monotonically damping entropy. The second is a double vortex where a constrained fully discrete formulation is shown to achieve exact entropy conservation in time.
@@ 1. Introduction
Mathematical entropies, or entropy, are convex functionals arising from a positive definite Hessian [Fisher2013, Leveque1992, Chapter 3] that correspond to quadratic buoyancy–related tracer invariants for the thermal shallow water equations [Ricardo2023, Ricardo2024-dg]. The thermal shallow water equations are a useful stepping stone from simpler atmospheric systems, like the rotating shallow water equations, to the full three-dimensional compressible Euler equations typically used to describe atmospheric dynamics in operational weather models [Maynard2020, Adams2019, Lee2021]. The thermal shallow equations are analogous to the compressible Euler equations with an identical Poisson bracket and entropy, and include a thermodynamic scalar quantity reflective of temperature [Eldred2019, Ricardo2023, Ricardo2024-dg]. Such thermodynamic quantities are not included in the rotating shallow water equations [Cotter2014, BauerCotter2018, Mcrae2013, Wimmer2020], meaning these other models do not capture the effect of thermodynamic transport on the horizontal pressure gradient [Kurganov2021]. In this study, we develop a novel finite element approximation of the thermal shallow water equations that provably conserves energy and entropy at the semi-discrete level. Large scale simulations show the resulting scheme is conservative, captures turbulent dynamics, and can stably simulate non-linear flow over long time scales where a mature turbulent state is reached.
@@ 2.2. System invariants
To motivate the discrete approximation proposed in Sections 3 and 4 below, we discuss entropy conservation at the continuous level. Hereafter, we drop the dependence on space and time. Assuming $b,\varphi,B\in H^{1}(J)$, the total time derivative of $\mathcal{S}$ is
$\frac{\mathrm{d}\mathcal{S}}{\mathrm{d}t}=\int_{\Omega}b\varphi\frac{\partial b}{\partial t}+\int_{\Omega}\frac{1}{2}b^{2}\frac{\partial\varphi}{\partial t}=\int_{\Omega}\left(b\frac{\partial B}{\partial t}-b^{2}\frac{\partial\varphi}{\partial t}\right)+\int_{\Omega}\frac{1}{2}b^{2}\frac{\partial\varphi}{\partial t}$ $=\int_{\Omega}\frac{\delta\mathcal{S}}{\delta B}\frac{\partial B}{\partial t}+\int_{\Omega}\frac{\delta\mathcal{S}}{\delta\varphi}\frac{\partial\varphi}{\partial t},$ (5)
@@
where $\delta\mathcal{S}/\delta B=b$ and $\delta\mathcal{S}/\delta\varphi=-b^{2}/2$ are the variational derivatives of entropy. Next, substitute 2b and 2c into 5. Expanding spatial derivatives and assuming $\nabla(b^{2})\in H^{1}(\Omega)$, $\boldsymbol{F}\in H(\mathrm{div},\Omega)$ in order to apply integration by parts gives
$\frac{\mathrm{d}\mathcal{S}}{\mathrm{d}t}$ $=\int_{\Omega}\frac{1}{2}b^{2}\nabla\cdot\boldsymbol{F}-\int_{\Omega}b\nabla\cdot(b\boldsymbol{F})=\int_{\Omega}\frac{1}{2}\nabla(b^{2})\cdot\boldsymbol{F}-\int_{\Omega}b\boldsymbol{F}\cdot\nabla b=0.$ (6)
@@
The above analysis is used to determine the regularity requirements for discrete entropy conservation. For 5 to hold discretely in time, we consider a temporally continuous approximation of $b$ and the prognostic variables. Thus, entropy is continuous cubic polynomial in time, which is not conserved in general at the discrete level by the chosen time integrator [Cohen2011]. The spatial approximation of prognostic variables is determined by a discrete de Rham complex [Cotter2023], where $\nabla\cdot\boldsymbol{u},\varphi,B$ are spatially discontinuous. For the vector invariant form of the thermal shallow water equations, $b$ is typically also spatially discontinuous [Ricardo2023, Eldred2019]. This choice means entropy conservation in 6 does not hold at the semi-discrete level [Ricardo2023]. While the material form of the thermal shallow water equations facilitates spatially continuous approximations of $b$, the associated variational derivative of the Hamiltonian is not collocated with the pressure gradient, and energy is not readily conserved [Eldred2019].
@@
To allow for spatially discontinuous approximations of $b$, we define a numerical approximation of 2 such that integration by parts is not required to prove semi-discrete entropy conservation. Using the chain rule to reformulate the buoyancy equation 2c and reciprocating modifications in the momentum equation 2a, we restate 2 as [Ricardo2023]
@@
Skew symmetry is maintained in 7 so that energy is conserved, assuming periodic spatial boundary conditions. Substituting 7b and 7c into 5, yields entropy conservation at the continuous level as in 6. At the discrete level, we apply integration by parts to $\nabla\cdot(b\boldsymbol{F})$ in 7c and account for jumps across interior element boundaries such that the resulting term readily cancels with the jumps related to the $\boldsymbol{F}\cdot\nabla b$ term in 7c. The $b\nabla\cdot\boldsymbol{F}$ term in 7c is div-conforming at the discrete level, and does not contribute to interior element boundaries. To ensure discrete energy and entropy conservation, we design skew symmetric jump terms by exploiting the regularity of $b$ at the continuous level.
@@ 3. Semi-discrete formulation
We now present a compatible finite element formulation of the thermal shallow water equations 7, and prove semi-discrete energy and entropy conservation. Skew symmetric upwinded numerical fluxes, which provably conserve energy and monotonically damp entropy, are proposed to smooth spurious oscillations that can occur as simulations progress to mature turbulent states.
@@ 3.1. Notation
$\llbracket\boldsymbol{y}\rrbracket$ $=\boldsymbol{y}^{+}\boldsymbol{n}^{+}+\boldsymbol{y}^{-}\boldsymbol{n}^{-},$$\left\{\boldsymbol{y}\right\}$ $=\frac{1}{2}\left(\boldsymbol{y}^{+}+\boldsymbol{y}^{-}\right)\text{ on }e\in\mathcal{E}_{0},$
@@ 3.3. Finite element approximation
The semi-discrete finite element formulation of the thermal shallow water equations in 7 is: find $\boldsymbol{u}_{h}\in\mathbb{V}_{1}$, $\varphi_{h},B_{h}\in\mathbb{V}_{2}$, such that
@@
for all $\boldsymbol{w}_{h}\in\mathbb{V}_{1}$, $\phi_{h}\in\mathbb{V}_{2}$, where $\widetilde{b_{h}}$ is a specific formulation of the buoyancy introduced in 25 below, and $\alpha(\boldsymbol{w}_{h})\coloneqq|\boldsymbol{w}_{h}\cdot\boldsymbol{n}^{+}|/2$ is the upwinding parameter [Brezzi2004]. Centred fluxes correspond to $\alpha=0$ and upwinded fluxes correspond to $\alpha>0$. The upwinded fluxes in 9c align with other studies that consider hyperbolic systems involving advection [Brezzi2004, Eldred2019], and maintain skew symmetry in 8. Upwinding is not required in the continuity equation since the corresponding linearised equation is balanced with respect to left and right solutions [Vallis2006], and $\boldsymbol{F}_{h}\in\mathbb{V}_{1}$ is continuous across element boundaries.
@@
$\left(\Phi_{h},\phi_{\Phi_{h}}\right)$ $=\left(\frac{1}{2}\boldsymbol{u}_{h}\cdot\boldsymbol{u}_{h}+B_{h},\phi_{\Phi_{h}}\right),$ $\forall\phi_{\Phi_{h}}\in\mathbb{V}_{2},$ (10b)
@@
where 10c holds point-wise and is not required in practice. The final diagnostic is: find $\widetilde{b_{h}}\in\mathbb{V}_{2}$, such that
@@
where the projection of the quadratic term $b_{h}b_{h}$ into $\mathbb{V}_{2}$ is accounted for since $\widetilde{\phi}_{h}\in\mathbb{V}_{2}$. While entropy can be conserved at the semi-discrete level by replacing $\widetilde{b_{h}}$ with $b_{h}$ [Ricardo2023, Ricardo2024-dg], we show in Section 4.3 that the additional diagnostic 11 ensures the forcing terms exactly conserve entropy in time. Thus, the only source of temporal conservation error is the approximation of time derivatives.
@@
To show smooth solutions of the continuous system satisfy the semi-discrete system, discrete quantities in 8 are replaced by their continuous analogue. Equation 8 contains terms involving the time derivative of the prognostic variables, for which the consistency is trivial, and all other terms relate to the diagnostic variables. We now elaborate on the consistency of the forms $g(\cdot,b_{h},\widetilde{b_{h}},\cdot)$ and $s(\cdot,b_{h},\cdot)$.
@@
after applying integration by parts and the chain rule. Consistency readily follows noting that $\nabla\cdot\psi_{\boldsymbol{u}}\Phi$ is div-conforming. We proceed analogously for the density weighted buoyancy equation 8c. ∎
@@
Setting $b_{h}=1$ in 10d and 11 yields $\varphi_{h}=B_{h}$ point-wise and $\widetilde{b_{h}}=1$ respectively. Next taking $b_{h}=\widetilde{b_{h}}=1$ in 8c, and expanding $g(\boldsymbol{F}_{h},1,1,\phi_{B_{h}})$ and $s(\boldsymbol{F}_{h},1,\phi_{B_{h}})$ gives
@@
since $\llbracket 1\rrbracket=\nabla_{h}1=0$ and $\left\{1\right\}=1$. Applying integration by parts yields 8b since $\varphi_{h}=B_{h}$. ∎
@@ 3.4. Semi-discrete conservation
We now prove the main result of this study: the semi-discrete system 8 conserves discrete entropy when $\alpha=0$, monotonically damps entropy when $\alpha>0$, and conserves energy always.
@@
$\frac{\delta\mathcal{H}_{h}}{\delta\varphi_{h}}\in\mathbb{V}_{2}:$ $\left(\frac{\delta\mathcal{H}_{h}}{\delta\varphi_{h}},\phi_{\varphi_{h}}\right)$ $=\left(\frac{1}{2}\boldsymbol{u}_{h}\cdot\boldsymbol{u}_{h}+B_{h},\phi_{\varphi_{h}}\right)=\left(\Phi_{h},\phi_{\varphi_{h}}\right),$ $\forall\phi_{\varphi_{h}}\in\mathbb{V}_{2},$ (15b)
@@
are the functional derivatives of $\mathcal{H}_{h}(t)$ that hold point-wise. Substituting 15 into 14 and evaluating the resulting expression using $\psi_{\boldsymbol{u}_{h}}=\boldsymbol{F}_{h}$, $\phi_{\varphi_{h}}=\Phi_{h}$, $\phi_{B_{h}}=\vartheta_{h}$ in 8 gives ${\partial\mathcal{H}_{h}}/{\partial t}=0$. ∎
@@
Under the assumption of continuity of time, 5 gives the time derivative of $\mathcal{S}_{h}$ as
@@
In contrast to energy, entropy involves both prognostic and diagnostic variables. Evaluating the discrete functional derivatives requires several steps as follows. First differentiating 16 with respect to $\varphi_{h}$ and $b_{h}$ gives
@@
Further differentiating 10d with respect to $\varphi_{h}$ and $B_{h}$ yields
@@
Thus, the variational derivatives of $\mathcal{S}_{h}(t)$ are:
@@
In obtaining 20a, we set $\phi_{b_{h}}={\partial b_{h}}/{\partial\varphi_{h}}$ in 18b, $\phi_{\varphi_{h}}=b_{h}$ in 19a, and combine the result with 18a. Deriving 20b requires taking $\phi_{b_{h}}={\partial b_{h}}/{\partial B_{h}}$ in 18b and $\phi_{b_{h}}=b_{h}$ in 19b.
To evaluate 17, set $\phi_{\varphi_{h}}=-\Pi_{\mathbb{V}_{2}}(b_{h}b_{h})/2$ in 8b and $\phi_{B_{h}}=b_{h}$ in 8c, where $\Pi_{\mathbb{V}_{2}}$ is the projection into $\mathbb{V}_{2}$. Expanding $g(\boldsymbol{F}_{h},b_{h},\widetilde{b_{h}},b_{h})$ and $s(\boldsymbol{F}_{h},b_{h},b_{h})$ in the resulting expression shows integration by parts is not required to cancel terms. Thus,
@@
where the projection into $\mathbb{V}_{2}$ cancels since $\nabla\cdot\boldsymbol{F}_{h}\in\mathbb{V}_{2}$. Using $\widetilde{\phi}_{h}=\nabla\cdot\boldsymbol{F}_{h}$ in 11 yields ${\partial\mathcal{S}_{h}}/{\partial t}=0$ when $\alpha=0$, and ${\partial\mathcal{S}_{h}}/{\partial t}<0$ when $\alpha>0$ since $s_{\mathrm{up}}(\boldsymbol{F}_{h},{b_{h}},{b_{h}})>0$ for $\alpha>0$. ∎
@@
Discrete mass $\mathcal{M}_{h}(t)=\int_{\Omega}\varphi_{h}$ is conserved since setting $\phi_{\varphi_{h}}=1$ in 8b results in a flux form conservation law. Taking $\phi_{B_{h}}=1$ in 8c and applying integration by parts reveals the discrete total buoyancy, $\mathcal{B}_{h}=\int_{\Omega}B_{h}$, is only conserved when $\widetilde{b_{h}}=b_{h}$ point-wise. Similar to other finite element approximations of the thermal shallow water equations [Eldred2019], discrete vorticity, $\mathcal{V}_{h}=\int_{\Omega}\varphi_{h}q_{h}$, is conserved when $b_{h}$ is a constant. Thus, the semi-discrete formulation satisfies the criteria of a compatible spatial discretisation for the thermal shallow water equations [Eldred2019, Ricardo2024-dg].

### v2

@@ Abstract.
In this work, we develop a new compatible finite element formulation of the thermal shallow water equations that conserves energy and mathematical entropies given by buoyancy–related quadratic tracer variances. Our approach relies on restating the governing equations to enable discontinuous approximations of thermodynamic variables and a variational continuous time integration. A key novelty is the inclusion of centred and upwinded fluxes. The proposed semi-discrete system conserves discrete entropy for centred fluxes, monotonically damps entropy for upwinded fluxes, and conserves energy. The fully discrete scheme preserves entropy conservation at the continuous level. The ability of a new linearised Jacobian, which accounts for both centred and upwinded fluxes, to capture large variations in buoyancy and simulate thermally unstable flows for long periods of time is demonstrated for two different transient case studies. The first involves a thermogeostrophic instability where including upwinded fluxes is shown to suppress spurious oscillations while successfully conserving energy and monotonically damping entropy. The second is a double vortex where a constrained fully discrete formulation is shown to achieve exact entropy conservation in time.
@@ 1. Introduction
Mathematical entropies, or entropy, are convex functionals arising from a positive definite Hessian [Fisher2013, Leveque1992, Chapter 3] that correspond to quadratic buoyancy–related tracer invariants for the thermal shallow water equations [Ricardo2023, Ricardo2024-dg]. The thermal shallow water equations are a useful stepping stone from simpler atmospheric systems, like the rotating shallow water equations, to the full three-dimensional compressible Euler equations typically used to describe atmospheric dynamics in operational weather models [Maynard2020, Adams2019, Lee2021]. The thermal shallow equations are analogous to the compressible Euler equations with an identical Poisson bracket and entropy, and include a thermodynamic scalar quantity reflective of temperature [Eldred2019, Ricardo2023, Ricardo2024-dg]. Such thermodynamic quantities are not included in the rotating shallow water equations [Cotter2014, BauerCotter2018, Mcrae2013, Wimmer2020], meaning this specific model does not capture the effect of thermodynamic transport on the horizontal pressure gradient [Kurganov2021]. In this study, we develop a novel finite element approximation of the thermal shallow water equations, and prove conservation of energy and entropy at the semi-discrete level. Large scale simulations show the resulting scheme is conservative, captures turbulent dynamics, and can stably simulate non-linear flow over long time scales where a mature turbulent state is reached.
@@ 2.2. System invariants
To motivate the discrete approximation proposed in Sections 3 and 4 below, we discuss entropy conservation at the continuous level. Hereafter, we drop the dependence on space and time. Assuming $b,\varphi,B\in H^{1}(J)$, differentiating 4 with respect to time yields the total time derivative of $\mathcal{S}$ as
$\frac{\mathrm{d}\mathcal{S}}{\mathrm{d}t}$ $=\int_{\Omega}b\varphi\frac{\partial b}{\partial t}+\int_{\Omega}\frac{1}{2}b^{2}\frac{\partial\varphi}{\partial t},$ (5a)
$=\int_{\Omega}B\frac{\partial}{\partial t}\left(B\varphi^{-1}\right)+\int_{\Omega}\frac{1}{2}b^{2}\frac{\partial\varphi}{\partial t},$ (5b)
$=\int_{\Omega}B\varphi^{-1}\frac{\partial B}{\partial t}-\int_{\Omega}B^{2}\varphi^{-2}\frac{\partial\varphi}{\partial t}+\int_{\Omega}\frac{1}{2}b^{2}\frac{\partial\varphi}{\partial t},$ (5c)
$=\int_{\Omega}b\frac{\partial B}{\partial t}-\int_{\Omega}\frac{1}{2}b^{2}\frac{\partial\varphi}{\partial t}.$ (5d)
To obtain 5b, we substitute the relation $b=B\varphi^{-1}$ into the first term of 5a. Expanding the temporal derivatives in the first term of 5b gives 5c. Further substituting $b=B\varphi^{-1}$ into the first and second terms of 5c and simplifying yields 5d. An equivalent expression for 5d is
$\frac{\mathrm{d}\mathcal{S}}{\mathrm{d}t}=\int_{\Omega}\frac{\delta\mathcal{S}}{\delta B}\frac{\partial B}{\partial t}+\int_{\Omega}\frac{\delta\mathcal{S}}{\delta\varphi}\frac{\partial\varphi}{\partial t},$ (6)
where $\delta\mathcal{S}/\delta B=b$ and $\delta\mathcal{S}/\delta\varphi=-b^{2}/2$ are the variational derivatives of entropy.
Next, substituting 2b and 2c into 5d yields
$\frac{\mathrm{d}\mathcal{S}}{\mathrm{d}t}$ $=\int_{\Omega}\frac{1}{2}b^{2}\nabla\cdot\boldsymbol{F}-\int_{\Omega}b\nabla\cdot(b\boldsymbol{F}),$ (7a)
$=\int_{\Omega}\frac{1}{2}b^{2}\nabla\cdot\boldsymbol{F}-\int_{\Omega}b\left(b\nabla\cdot\boldsymbol{F}+\boldsymbol{F}\cdot\nabla b\right),$ (7b)
$=-\int_{\Omega}\frac{1}{2}b^{2}\nabla\cdot\boldsymbol{F}-\int_{\Omega}b\boldsymbol{F}\cdot\nabla b,$ (7c)
@@
where 7b results from expanding the spatial derivatives in the second term of 7a, and 7c arises from grouping terms. Now we assume $\nabla(b^{2})\in H^{1}(\Omega)$, $\boldsymbol{F}\in H(\mathrm{div},\Omega)$ in order to apply integration by parts to the first term in 7c. Expanding spatial derivatives in the resulting term gives
$\frac{\mathrm{d}\mathcal{S}}{\mathrm{d}t}$ $=\int_{\Omega}\frac{1}{2}\nabla(b^{2})\cdot\boldsymbol{F}-\int_{\Omega}b\boldsymbol{F}\cdot\nabla b=\int_{\Omega}b\boldsymbol{F}\cdot\nabla b-\int_{\Omega}b\boldsymbol{F}\cdot\nabla b=0.$ (8)
@@
The above analysis is used to determine the regularity requirements for discrete entropy conservation. For 6 to hold discretely in time, we consider a temporally continuous approximation of $b$ and the prognostic variables. Thus, entropy is a continuous cubic polynomial in time, which is not conserved in general at the discrete level by the chosen time integrator [Cohen2011]. The spatial approximation of prognostic variables is determined by a discrete de Rham complex [Cotter2023], where $\nabla\cdot\boldsymbol{u},\varphi,B$ are spatially discontinuous. For the vector invariant form of the thermal shallow water equations, $b$ is typically also spatially discontinuous [Ricardo2023, Eldred2019]. This choice means entropy conservation in 8 does not hold at the semi-discrete level [Ricardo2023]. While the material form of the thermal shallow water equations facilitates spatially continuous approximations of $b$, the associated variational derivative of the Hamiltonian is not collocated with the pressure gradient, and energy is not readily conserved [Eldred2019].
@@
To allow for spatially discontinuous approximations of $b$, we define a numerical approximation of 2 such that integration by parts is not required to prove semi-discrete entropy conservation. Using the product rule to reformulate the buoyancy equation 2c and reciprocating modifications in the momentum equation 2a, we restate 2 as [Ricardo2023]
@@
Skew-symmetry is maintained in 9 so that energy is conserved, assuming periodic spatial boundary conditions. Substituting 9b and 9c into 5d, yields entropy conservation at the continuous level as in 8. At the discrete level, we apply integration by parts to $\nabla\cdot(b\boldsymbol{F})$ in 9c and account for jumps across interior element boundaries such that the resulting term readily cancels with the jumps related to the $\boldsymbol{F}\cdot\nabla b$ term in 9c. The $b\nabla\cdot\boldsymbol{F}$ term in 9c is div-conforming at the discrete level, and does not contribute to interior element boundaries. To ensure discrete energy and entropy conservation, we design skew symmetric jump terms by exploiting the regularity of $b$ at the continuous level.
@@ 3. Semi-discrete formulation
We now present a compatible finite element formulation of the thermal shallow water equations 9, and prove semi-discrete energy and entropy conservation. Skew symmetric upwinded numerical fluxes are proposed to smooth spurious oscillations that can occur as simulations progress to mature turbulent states. We subsequently prove such fluxes conserve energy and monotonically damp entropy.
@@ 3.1. Notation
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\llbracket\boldsymbol{y}\rrbracket}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{y}^{+}\cdot\boldsymbol{n}^{+}+\boldsymbol{y}^{-}\cdot\boldsymbol{n}^{-}},$$\left\{\boldsymbol{y}\right\}$ $=\frac{1}{2}\left(\boldsymbol{y}^{+}+\boldsymbol{y}^{-}\right)\text{ on }e\in\mathcal{E}_{0},$
@@ 3.3. Finite element approximation
The semi-discrete finite element formulation of the thermal shallow water equations in 9 is: find $\boldsymbol{u}_{h}\in\mathbb{V}_{1}$, $\varphi_{h},B_{h}\in\mathbb{V}_{2}$, such that
@@
for all $\boldsymbol{w}_{h}\in\mathbb{V}_{1}$, $\phi_{h}\in\mathbb{V}_{2}$, where $\widetilde{b_{h}}$ is a specific formulation of the buoyancy introduced in 27 below, and $\alpha(\boldsymbol{w}_{h})\coloneqq|\boldsymbol{w}_{h}\cdot\boldsymbol{n}^{+}|/2$ is the upwinding parameter [Brezzi2004]. Centred fluxes correspond to $\alpha=0$ and upwinded fluxes correspond to $\alpha>0$. The upwinded fluxes in 11c align with other studies that consider hyperbolic systems involving advection [Brezzi2004, Eldred2019], and maintain skew symmetry in 10. Upwinding is not required in the continuity equation since the corresponding linearised equation is balanced with respect to left and right solutions [Vallis2006], and $\boldsymbol{F}_{h}\in\mathbb{V}_{1}$ is continuous across element boundaries.
@@
$\left(\Phi_{h},\phi_{\Phi_{h}}\right)$ $=\left(\frac{1}{2}\boldsymbol{u}_{h}\cdot\boldsymbol{u}_{h}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{2}}B_{h},\phi_{\Phi_{h}}\right),$ $\forall\phi_{\Phi_{h}}\in\mathbb{V}_{2},$ (12b)
@@
where 12c holds pointwise and is not required in practice. The final diagnostic is: find $\widetilde{b_{h}}\in\mathbb{V}_{2}$, such that
@@
where the projection of the quadratic term $b_{h}b_{h}$ into $\mathbb{V}_{2}$ is accounted for since $\widetilde{\phi}_{h}\in\mathbb{V}_{2}$. While entropy can be conserved at the semi-discrete level by replacing $\widetilde{b_{h}}$ with $b_{h}$ [Ricardo2023, Ricardo2024-dg], we show in Section 4.3 that the additional diagnostic 13 ensures the forcing terms exactly conserve entropy in time. Thus, the only source of temporal conservation error is the approximation of time derivatives.
@@
To show smooth solutions of the continuous system satisfy the semi-discrete system, discrete quantities in 10 are replaced by their continuous analogue. Equation 10 contains terms involving the time derivative of the prognostic variables, for which the consistency is trivial, and all other terms relate to the diagnostic variables. We now elaborate on the consistency of the forms $g(\cdot,b_{h},\widetilde{b_{h}},\cdot)$ and $s(\cdot,b_{h},\cdot)$.
@@
after applying integration by parts and the chain rule. Consistency readily follows noting that $\nabla\cdot\psi_{\boldsymbol{u}}\Phi$ is div-conforming. A similar procedure applies to the density weighted buoyancy equation 10c.
@@
Setting $b_{h}=1$ in 12d and 13 yields $\varphi_{h}=B_{h}$ pointwise and $\widetilde{b_{h}}=1$ respectively. Next taking $b_{h}=\widetilde{b_{h}}=1$ in 10c, and expanding $g(\boldsymbol{F}_{h},1,1,\phi_{B_{h}})$ and $s(\boldsymbol{F}_{h},1,\phi_{B_{h}})$ gives
@@
since $\llbracket 1\rrbracket=\nabla_{h}1=0$ and $\left\{1\right\}=1$. Applying integration by parts yields 10b since $\varphi_{h}=B_{h}$. ∎
@@ 3.4. Semi-discrete conservation
We now prove the main result of this study: the semi-discrete system 10 conserves discrete entropy when $\alpha=0$, monotonically damps entropy when $\alpha>0$, and conserves energy always.
@@
$\frac{\delta\mathcal{H}_{h}}{\delta\varphi_{h}}\in\mathbb{V}_{2}:$ $\left(\frac{\delta\mathcal{H}_{h}}{\delta\varphi_{h}},\phi_{\varphi_{h}}\right)$ $=\left(\frac{1}{2}\boldsymbol{u}_{h}\cdot\boldsymbol{u}_{h}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{2}}B_{h},\phi_{\varphi_{h}}\right)=\left(\Phi_{h},\phi_{\varphi_{h}}\right),$ $\forall\phi_{\varphi_{h}}\in\mathbb{V}_{2},$ (17b)
@@
are the functional derivatives of $\mathcal{H}_{h}(t)$ that hold pointwise. Substituting 17 into 16 and evaluating the resulting expression using $\psi_{\boldsymbol{u}_{h}}=\boldsymbol{F}_{h}$, $\phi_{\varphi_{h}}=\Phi_{h}$, $\phi_{B_{h}}=\vartheta_{h}$ in 10 gives ${\partial\mathcal{H}_{h}}/{\partial t}=0$. ∎
@@
Under the assumption of continuity in time, 6 gives the time derivative of $\mathcal{S}_{h}$ as
@@
In contrast to energy, entropy involves both prognostic and diagnostic variables. Evaluating the discrete functional derivatives requires several steps as follows. First differentiating 18 with respect to $\varphi_{h}$ and $b_{h}$ gives
@@
Further differentiating 12d with respect to $\varphi_{h}$ and $B_{h}$ yields
@@
To obtain an expression for $\delta\mathcal{S}_{h}/\delta\varphi_{h}$, we set $\phi_{b_{h}}={\partial b_{h}}/{\partial\varphi_{h}}$ in 20b, $\phi_{\varphi_{h}}=b_{h}$ in 21a, and combine the result with 20a. Deriving an expression for $\delta\mathcal{S}_{h}/\delta B_{h}$ requires taking $\phi_{b_{h}}={\partial b_{h}}/{\partial B_{h}}$ in 20b and $\phi_{b_{h}}=b_{h}$ in 21b. Thus, the variational derivatives of $\mathcal{S}_{h}(t)$ are:
@@
To evaluate 19, set $\phi_{\varphi_{h}}=-\Pi_{\mathbb{V}_{2}}(b_{h}b_{h})/2$ in 10b and $\phi_{B_{h}}=b_{h}$ in 10c, where $\Pi_{\mathbb{V}_{2}}$ is the projection into $\mathbb{V}_{2}$. Expanding $g(\boldsymbol{F}_{h},b_{h},\widetilde{b_{h}},b_{h})$ and $s(\boldsymbol{F}_{h},b_{h},b_{h})$ in the resulting expression shows integration by parts is not required to cancel terms. Thus,
@@
where the projection into $\mathbb{V}_{2}$ cancels since $\nabla\cdot\boldsymbol{F}_{h}\in\mathbb{V}_{2}$. Using $\widetilde{\phi}_{h}=\nabla\cdot\boldsymbol{F}_{h}$ in 13 yields ${\partial\mathcal{S}_{h}}/{\partial t}=0$ when $\alpha=0$, and ${\partial\mathcal{S}_{h}}/{\partial t}<0$ when $\alpha>0$ since $s_{\mathrm{up}}(\boldsymbol{F}_{h},{b_{h}},{b_{h}})>0$ for $\alpha>0$. ∎
@@
Discrete mass $\mathcal{M}_{h}(t)=\int_{\Omega}\varphi_{h}$ is conserved since setting $\phi_{\varphi_{h}}=1$ in 10b results in a flux form conservation law. Taking $\phi_{B_{h}}=1$ in 10c and applying integration by parts reveals the discrete total buoyancy, $\mathcal{B}_{h}=\int_{\Omega}B_{h}$, is only conserved when $\widetilde{b_{h}}=b_{h}$ pointwise. Similar to other finite element approximations of the thermal shallow water equations [Eldred2019], discrete vorticity, $\mathcal{V}_{h}=\int_{\Omega}\varphi_{h}q_{h}$, is conserved when $b_{h}$ is a constant. Thus, the semi-discrete formulation satisfies the criteria of a compatible spatial discretisation for the thermal shallow water equations [Eldred2019, Ricardo2024-dg].